prove that the Kepler’s law of area satisfies the conservation principle of energy. 



1
Expert's answer
2021-03-08T08:30:14-0500

 Kepler’s law of area:


"\frac{dA}{dt}=\frac{r^2\omega}{2}=const"


"\frac{Mr^2\omega}{2M}=\frac{L}{2M}=const\\L=const"